I am very confused solve and round to nearest1000th
6z^2+5z=-1

Answer by stanbon(75887)

solve and round to nearest1000th
6z^2+5z=-1
---
6z^2 + 5z - 1 = 0
(6z -1)(z+1) = 0
z = 1/6 or z = -1
=======================
Cheers,
Stan H.